# Collective behavior of independent scaled Brownian particles with renewal resetting

### v1

@@ Abstract
[inst1] organization=Racah Institute of Physics, Hebrew University of Jerusalem, addressline=, city=Jerusalem, postcode=91904, state=, country=Israel
@@ 4 Full statistics of the center of mass
Then, by the law of large numbers, the variance of the center of mass $X$ is
@@ 4.1 $H\leq 1/2$: standard scaling
The starting point of the calculation is the exact one-particle position distribution $p_{\text{s}}(x)$ [Bodrova2019], described by Eq. (2). The joint steady-state distribution $P(x_{1},x_{2},\dots,x_{N})$ of the positions of all particles can be obtained from the single-particle distribution:
@@
Equations (33) and (34) determine the rate function $f(a;H)$ in a parametric form, where $\kappa$ plays the role of the parameter. Figure 4 shows plots of this rate function for $H=1/8$, $1/3$ and $1/2$.
@@ 5 Discussion
Acknowledgment. B. M. was supported by the Israel Science Foundation (Grant No. 1579/25).

### v2

@@ Abstract
[inst1] organization=Racah Institute of Physics, Hebrew University of Jerusalem, city=Jerusalem, postcode=91904, country=Israel
@@ 4 Full statistics of the center of mass
By virtue of additivity of the variance of independent random variables, the variance of the center of mass $X$ is
@@ 4.1 $H\leq 1/2$: standard scaling
In this case the rate function $f(a)$ is given by Cramér’s theorem, see e.g. Ref. [Petrov]. Still, for completeness, we will calculate $f(a)$ from first principles. The starting point of this calculation is the exact one-particle position distribution $p_{\text{s}}(x)$ [Bodrova2019], described by Eq. (2). The joint steady-state distribution $P(x_{1},x_{2},\dots,x_{N})$ of the positions of all particles can be obtained from the single-particle distribution:
@@
This expression reproduces Cramér’s theorem, see e.g. Ref. [Petrov]. Equations (33) and (34) determine the rate function $f(a;H)$ in a parametric form, where $\kappa$ plays the role of the parameter. Figure 4 shows plots of this rate function for $H=1/8$, $1/3$ and $1/2$.
@@ 5 Discussion
Acknowledgments. We are grateful to Naftali R. Smith for valuable comments. B. M. was supported by the Israel Science Foundation (Grant No. 1579/25).